# Kernel Three Pass Regression Filter

### v1

@@ 5.7 Forecasting Housing Variables
Results and Discussions–
Table-9 presents the forecasting results. One can note that predicting the privately owned housing start time series is very difficult. While most forecasting methods cannot beat the historical average, our method performs better than all other methods at all horizons. On the other hand, while it is relatively easy to forecast housing in the western census region, our method performs better than all other methods except for one or two cases. The results in other housing variables follow similar patterns; therefore, we omit them for the simplicity of exposition.
Therefore, it is clear that our method holds value in a difficult forecasting problem in both the short and long run. We also enjoy a better method for long-run forecasting problems in the housing market series.
5.8 Forecasting Price Variables
Inflation, a topic of significant interest in the market, media, and academia, holds a crucial role due to its impact on the common consumers. Particularly, inflation adversely affects low-wage earners, underscoring its societal implications. The relevance of our research is underscored by the fact that one of the primary objectives of the Federal Reserve or Central Bank is to control inflation. In this context, we focus on forecasting two key price time series: the GDP Deflator and the Consumer Price Index (CPI), which are pivotal in understanding and predicting inflation. The motivation behind choosing these two price series is that– on the one hand, the GDP Deflator considers the macroeconomic picture at the aggregate level. At the same time, the CPI measure informs us of the inflation faced by the consumer at a disaggregated level.
GDP Deflator
Method h=1 h=2 h=3 h=4 h=6 h=8 h=10 h=12
AR(2) 0.797 0.774 0.740 0.657 0.418 0.146 0.156 0.077
PCA 0.444 0.276 0.056 -0.184 -0.408 -0.347 -0.221 -0.057
Sq-PC 0.299 0.145 -0.035 -0.168 -0.245 -0.230 -0.192 -0.108
PC-Sq 0.431 0.268 0.104 -0.039 -0.106 -0.038 -0.111 -0.182
kPCA -0.032 0.247 -0.021 0.008 0.003 0.004 0.029 -0.023
3PRF 0.584 0.496 0.426 0.243 0.174 0.279 0.300 0.155
k3PRF 0.667 0.632 0.563 0.476 0.479 0.413 0.197 0.512
CPI
Method h=1 h=2 h=3 h=4 h=6 h=8 h=10 h=12
AR(2) 0.704 0.706 0.620 0.565 0.397 0.211 0.062 -0.038
PCA 0.660 0.535 0.364 0.154 -0.163 -0.252 -0.248 -0.173
Sq-PC 0.410 0.296 0.161 0.049 -0.055 -0.156 -0.200 -0.173
PC-Sq 0.649 0.512 0.353 0.186 -0.019 -0.087 -0.187 -0.228
kPCA 0.440 0.380 -0.050 0.189 -0.043 -0.024 0.042 -0.006
3PRF 0.641 0.566 0.487 0.352 0.192 0.241 0.255 0.141
k3PRF 0.676 0.612 0.541 0.463 0.469 0.434 0.349 0.477
Table 10: Out of Sample $R^{2}$ of Price Variables (Chain-type Price Indices)
Results and Discussion
Table-10 presents the results. It reiterates almost the same story we discussed in previous subsections. In both series, we dominate all other methods at all forecast horizons. While we marginally lag the AR model at some short horizons, we decisively beat it in longer horizons.
5.9 Forecasting Financial Variables
@@
Table 11: Out of Sample $R^{2}$ of Financial Market Series
While the necessity for forecasting financial variables may seem self-evident, it is pertinent to underscore the significance of such endeavors. The collective market capitalization of companies constituting the S&P 500 stands at approximately 43 trillion US dollars, underscoring the substantial economic weight of these entities. While stock market indices, including the S&P 500, are typically considered risky financial assets, government securities represent their low-risk counterpart. Within this sub-section, we forecast both risky (S&P 500 Index), safe (GS-1), and one in-between (GS-10) financial series.
Results and Discussion
There is nothing new regarding the patterns of the forecasting performance we are observing. The GS-1 means government securities (Treasury Bills), which mature in one year, and similarly, the GS-10 means the treasury notes, which mature in 10 years. Our method beats all other methods except the AR model on all forecast horizons for all the series discussed here. While marginally lag the AR model in the short run, we beat it decisively in the longer horizon forecasts.
6 Comparing the Performance on All Series
To enhance the robustness of our empirical analysis, we conducted comparative assessments of our method against competing methods across all 176 series within our dataset. This entailed selecting each series as the target and repeating the comparative analysis for every series in our dataset. This subsection reports comprehensive competitive assessments of forecasting that are under consideration against our method.
@@ 6.1 Description of Comparisons
We conducted the forecast comparisons across eight horizons, denoted as $h={1,2,3,4,6,8,10,12}$. Consequently, our investigation encompasses the comparative performance of models across a total of $176\times 8=1408$ combinations. The results of these comparisons, indicating the percentage of instances where a particular method demonstrated superior performance, are presented in Table 17. For example, if a method emerged as the best performer in 704 out of 1408 combinations, it would be represented by a value of 50 in the table. In simpler words, we count the relative frequency of the occurrence of the best performance of a given method and express it in percentage terms.
While the preceding frequency comparisons provide insight into the number of times each method proved superior to others, they do not measure the extent to which the best-performing method surpassed its nearest competitor. In other words, while method A may marginally outperform method B on one forecast horizon, method B might exhibit a considerable advantage over method A on another horizon, and then the aforementioned frequency comparison may not present the full picture. To account for this measure, we introduce a notion of ‘Tolerance’ level. We call a method ‘best’ under tolerance level $\epsilon$ if the out-of-sample $R^{2}$ of a method is within $\epsilon$ percentage lower than the best method’s performance. For example, suppose the AR model is the best for a combination of series $y$ and $h=h_{0}$ with a $R^{2}=0.60$. For tolerance=5, another method will also be considered ‘best’ if its $R^{2}$ is greater than or equal to $0.60(1-5/100)=0.57$. Therefore, for a non-zero tolerance, it is possible to have multiple ‘best’ methods. When we set tolerance=0, the method counts relative frequencies.
The “Overall” set of rows presents the percentage of instances where a method outperformed all other methods or lies in the tolerance range of the best-performing method among all 1408 comparisons. Recognizing that forecast objectives may vary in time horizon, we scrutinize comparative performances in short- and long-run contexts. The “Short-run” rows incorporate horizons $h={1,2,3,4}$, comprising 708 (calculated as $176\times 4$) combinations, while the “Long-run” row includes horizons $h={6,8,10,12}$, similarly amounting to 708 combinations. Additionally, we present a row labeled “Without AR”, wherein we exclude the auto-regressive method and compare the remaining methodologies. Going beyond short and long-run analysis, we enrich our analysis by reporting similar comparative performance numbers for each forecast horizon $h$. These numbers are reported in the appendix-B.5.
Before closing this section, we want to clarify that since multiple ‘best’ methods are possible for a non-zero tolerance level, the sum of rows can exceed 100 percent for non-zero tolerance. For tolerance=0, the rows sum up to 100 percent.
@@ 6.2 Results
We present the results into table-12. The findings presented in this study yield several noteworthy observations. First, it is evident that unsupervised forecasting techniques, including PCA, Squared-PC, PC-squared, and kernel PCA, exhibit inferior performance across the majority of scenarios when compared to our method. Second, our method, kernel 3PRF, demonstrates unequivocal superiority in long-term forecasting endeavors. Third, our method is unequivocally superior in the absence of the AR method. Our method does not outperform AR in the short term, but its performance remains competitive, often closely trailing the best short-run autoregressive method. This can be seen by increasing the tolerance level. We can see that the instances where our method can be labeled as ‘best’ increase rapidly as we increase the tolerance level.
In summary, the table offers insights into the overall efficacy of our method across various forecast horizons. While the autoregressive (AR) method may exhibit marginal superiority in short-term forecasts, this advantage is notably slim. Conversely, when our method surpasses AR, it does so considerably. Consequently, our method emerges as a dependable and preferred choice for forecasting across all horizons, offering robust and consistent performance.
@@ 7 Conclusion
Building upon the three-pass regression filter by Kelly & Pruitt, (2015), we introduce a new forecasting method, kernel three-pass regression filter. Our contention is that this approach holds promise as a dependable forecasting tool in environments characterized by approximate latent factor structure in predictors owing to two primary factors.
Firstly, it adeptly integrates non-linear relationships by transforming input data into a higher-dimensional space encapsulating its non-linear functions. Secondly, it operates as a supervised method, effectively filtering out and discarding useless factors to predict the target variable. We prove that the forecast generated by our method converges to the best infeasible forecast in probability as both $N$ and $T$ grow large.
We compare our method with competitor methods on 176 times series in FRED-QD data. Our findings consistently indicate that our methodology outperforms all competitors, particularly in longer-term forecasting scenarios across most time series under examination. The autoregressive (AR) method is the sole contender in short-term forecasting challenges. While the AR method may exhibit a marginal edge in short-term predictions, the differential advantage is negligible. Conversely, in instances where our methodology surpasses the AR method, it does so significantly. Consequently, our approach is dependable and preferred for forecasting across all temporal horizons, ensuring robust and steadfast performance.
@@ A.1 Proofs of Theoretical Results
The final line follows directly from Lemma 1.
@@
The final line follows directly from Lemma 2.
@@ A.2 Mercer’s Theorem
Appendix B Algorithms, Data, and Figures
@@
B.1 Out of Sample Estimation
@@ 4 Estimate $\boldsymbol{\beta}$ using the following formula:
We have demonstrated the construction of an out-of-sample forecast (Table-13). Now, we outline the rolling window procedure to obtain the out-of-sample forecast performance measured by out-of-sample $R^{2}$ in Table-14.
@@ 3 Compute Out-of-sample $R^{2}$:
B.2 Data Source and Description
We use FRED-QD data. This section provides the codes of the variables we forecast in our empirical work. For detailed description details, refer to FRED website. In table-15, FRED means federal reserve economic data, and SW stands for Stock and Watson datasets.
@@ 3 Decimal (Billions of Chained 2012 Dollars)
B.3 Hyper-parameter Tuning Algorithm
@@ 1. Take training input data $\{X_{train},y_{train}\}$ and split it into two halves:
B.4 Comparative Forecast Performance
We plot the forecasts by our method and 3PRF method with the true value of the target series for all sixteen series discussed in the emprirical application section. To save some space, we only show the plots for one, four, eight, and twelve period ahead forecasts.
@@ B.4 Comparative Forecast Performance
B.5 Comparative Performance on All Series For Each Horizons

### v3

@@ 5.6 Forecasting Housing and Financial Variables
Table 9: Out of Sample $R^{2}$ of Housing and Financial Variables
As seen from Table-9, forecasting HStart(total) proves to be a difficult problem. While most forecasting methods do not beat the historical average, our method performs better than all other methods at all horizons. It is relatively easy to forecast housing in the western census region, and our method performs better than all other methods except for a few cases. We find similar patterns in financial market variables, thereby omitting discussion.
6 Comprehensive Forecasting Analysis
To enhance the robustness of our empirical analysis, we conducted comparative assessments of our method against competing methods across all 176 series within our dataset. This entailed selecting each series as the target and repeating the comparative analysis for every series in our dataset.
@@ 6.1 Description of Comparisons
Our investigation encompasses the comparative performance of models across a total of $176\times 8=1408$ target-horizon combinations. The results of these comparisons, indicating the percentage of instances where a particular method demonstrated superior performance, are presented in Supplementary appendix-B.7. For example, if a method emerged as the best performer in 704 out of 1408 combinations, it would be represented by a value of 50 in the table. Essentially, we list the relative frequency of the occurrence of the best performance of a given method.
While the preceding frequency comparisons provide insight into the number of times each method proved superior to others, they do not measure the extent to which the best-performing method surpassed its nearest competitor. In other words, while method A may marginally outperform method B on one forecast horizon, method B might exhibit a considerable advantage over method A on another horizon. Then, the aforementioned frequency comparison may not depict the full picture. To account for this, we introduce a notion of ‘Tolerance’ level. We call a method ‘best’ under tolerance level $\epsilon$ if the out-of-sample $R^{2}$ of a method is within $\epsilon$ percentage lower than the best method’s performance  [^13] For example, if the AR model is the best for a of series $y_{\ell}$ and horizon $h_{0}$ with a $R^{2}=0.60$. For tolerance=5, another method will also be considered ‘best’ if its $R^{2}\geq 0.60(1-5/100)=0.57$. Therefore, for a non-zero tolerance, it is possible to have multiple ‘best’ methods.
The “All Horizons” set of rows summarizes all 1408 comparisons, i.e. encompassing all horizons and all series. Recognizing that forecast objectives may vary in time horizon, we scrutinize comparative performances in short- and long-run contexts. The “Short-run” rows incorporate horizons $h={1,2,3,4}$, comprising 708 (calculated as $176\times 4$) combinations, while the “Long-run” row includes horizons $h={6,8,10,12}$, similarly amounting to 708 combinations. Additionally, the portion labeled as “Excluding AR” excludes the auto-regressive method and compares the remaining methods across all 1408 combinations. For more granular analysis, we report comparative performance numbers for each forecast horizon $h$. These numbers are reported in the Supplementary appendix-B.7.
It is important to note that multiple ‘best’ methods may exist for a non-zero tolerance level, resulting in the sums of rows (in Table 5 in Supplementary appendix-B.7) exceeding 100 percent. However, for a tolerance level of zero, the rows sum to 100 percent.
@@ 6.2 Results
Table 10: Distribution of Best Forecasting Methods Across All Series (Percentage)
We present the results in table-10. The findings presented above yield several noteworthy observations. First, it is evident that unsupervised forecasting techniques, including PCR, Squared-PC, PC-squared, and kernel PCA, exhibit inferior performance across the majority of scenarios when compared to our method. Second, our method, kernel 3PRF, demonstrates unequivocal superiority in longer-horizon forecasting endeavors. Third, our method is unequivocally superior across all horizons when autoregressive (AR) method is excluded. Our method does not outperform AR in the short term, but its performance remains competitive, often closely trailing the best short-run autoregressive method. This can be seen by increasing the tolerance level. The instances where our method can be labeled as ‘best’ increase rapidly as we increase the tolerance level.
@@ 7 Conclusion
Building upon the three-pass regression filter by Kelly & Pruitt, (2015), we introduce a new forecasting method, kernel three-pass regression filter. Through extensive empirical exercises, we show that this approach holds promise as a dependable forecasting tool. Improved performance can be attributed to two noteworthy features of our method. First, it integrates non-linear relationships by transforming input data into a higher-dimensional space, encapsulating its non-linear functions. Second, it operates as a supervised method, effectively filtering out and discarding irrelevant factors while predicting the target variable.
@@ A.1 Proofs of Theoretical Results
The final line follows directly from Lemma 1 and Assumptions 2.1 and 2.2.
@@
The final line follows directly from Lemma 1 and Assumptions 2.1 and 2.2.
@@ A.2 Mercer’s Theorem
Appendix B Supplementary Appendix: Algorithms, Data, and Figures
@@
B.1 Some Popular Kernel Methods and Their Working
Polynomial Kernel
Then, the corresponding kernel function becomes:
$\mathcal{K}(a,b)=$ $\varphi(a)^{\prime}\varphi(b)$
Gaussian Kernel
This kernel is an example of an infinite-dimensional kernel. Let $\mathrm{x},\mathrm{z}\in\mathbb{R}^{k}$ and $\mathcal{K}(\mathrm{x},\mathrm{z})=e^{-\gamma\|\mathrm{x}-\mathrm{z}\|^{2}}$. Then, through the Taylor expansion, we can write
$=$ $\varphi(\mathrm{x})^{\prime}\varphi(\mathrm{z})$
B.2 Performance Metric: Out of Sample $R^{2}$
It is noteworthy that out-of-sample $R^{2}$ ranges from $\infty$- ∞ to $1$, unlike in-sample $R^{2}$, which ranges from zero to one. A positive out-of-sample $R^{2}$ indicates that the forecasting method outperforms the historical average. At the same time, a negative value suggests that the forecasting method performs worse than a simple method that forecasts $y_{i}$ equal to the historical average. We adopt the rolling window method to compute out-of-sample $R^{2}$, consistent with standard practices in the literature. Appendix B.3 provides a detailed exposition of our methodology.
B.3 Out of Sample Estimation
@@ 4 Estimate $\boldsymbol{\beta}$ using the following formula:
We have demonstrated the construction of an out-of-sample forecast (Table-11). Now, we outline the rolling window procedure to obtain the out-of-sample forecast performance measured by out-of-sample $R^{2}$ in Table-12.
@@ 3 Compute Out-of-sample $R^{2}$:
B.4 Data Source and Description
We use FRED-QD data. This section provides the codes of the variables we forecast in our empirical work. For detailed description details, refer to FRED website. In table-13, FRED means federal reserve economic data, and SW stands for Stock and Watson datasets.
@@ 3 Decimal (Billions of Chained 2012 Dollars)
B.5 Hyper-parameter Tuning Algorithm
@@ 1. Take training input data $\{X_{train},y_{train}\}$ and split it into two halves:
B.6 Comparative Forecast Performance
We plot the forecasts using our method and the 3PRF method with the true value of the target series for all sixteen series discussed in the empirical application section. To save some space, we only show the plots for one, four, eight, and twelve period ahead forecasts.
@@ B.6 Comparative Forecast Performance
B.7 Comparative Performance on All Series For Each Horizons